Find two numbers whose sum is 55 and product is 600.
Answer by oberobic(2304)

"Two numbers" suggests we use 'x' and 'y'.
However, we are told they sum to 55: x + y = 55
That means we can substitute 55-x for y.
That, in turn, means we have only one unknown to solve.
.
We also are told that x*y = 600
Substituting y = 55-x
.
x*(55-x) = 600
55x -x^2 = 600
.
Add x^2 to both sides
55x = x^2 + 600
.
Subtract 55x from both sides
0 = x^2 -55x + 600
.
Factor
.
(x-15)(x-40)=0
.
So we we have two possible values for x: 15 and 40.
That means we have two possible values for y:
55-15 = 40 and
55-40 = 15.
.
Answer:
The two numbers are 15 and 40. Their sum is 55. Their product is 600.
.
Done.